OpenStudy (anonymous):

if 2tanx/1-tan^2x=1, then x can equal____?

OpenStudy (anonymous):

guess you can start with \[2\tan(x)=1-\tan^2(x)\]

OpenStudy (anonymous):

or even \[\tan^2(x)+2\tan(x)-1=0\]

OpenStudy (anonymous):

\[tan(2x)= \dfrac{2tanx}{1-tan^2x}=1\\2x=arctan1\\2x = \dfrac{\pi}{4}\\x =\dfrac{\pi}{8}\]

OpenStudy (anonymous):

@mathmale can you help?

OpenStudy (shamim):

plz c my attachment